# ExBody2: Advanced Expressive Humanoid Whole-Body Control

### v1

@@ Abstract
This paper enables real-world humanoid robots to maintain stability while performing expressive motions like humans do. We propose Advanced Expressive Whole-Body Control (ExBody2), a generalized whole-body tracking framework that can take any reference motion inputs and control the humanoid to mimic the motion. The model is trained in simulation with Reinforcement Learning and then transferred to the real world. It decouples keypoint tracking with velocity control, and effectively leverages a privileged teacher policy to distill precise mimic skills into the target student policy, which enables high-fidelity replication of dynamic movements such as running, crouching, dancing, and other challenging motions. We present a comprehensive qualitative and quantitative analysis of crucial design factors in the paper. We conduct our experiments on two humanoid platforms and demonstrate the superiority of our approach against state-of-the-arts, providing practical guidelines to pursue the extreme of whole-body control for humanoid robots. More results can be found at https://exbody2.github.io.
@@
Figure 1: Humanoid robot executing various expressive whole-body motions in the real world. The robot can (a) walk with a large stride from static standing, (b): dance along a long horizon (43 seconds) choreography, (c) crouch and stand up, (d) punch with different height configurations, (e,f) express various upper-body movements while maintaining balance.
1 Introduction
Humanoid robots, with their similar morphology to humans, possess the potential to perform various tasks and motions that humans can accomplish in daily life. However, due to the high-dimensional state space and the complexity of control, it remains challenging to develop emergent human-like behaviors, limiting their real-world applications. With the growing availability of large-scale human motion datasets [4, 45], a practical approach to address this challenge is learning to replicate versatile movements by tracking and mimicking human motions [8, 23, 24, 20]. However, when considering the hardware, humanoid robots and humans are still quite different, which hinders the robot’s ability to replicate human motion completely. This poses a compelling research question: how can we pursue the expressive, human-like capabilities of humanoid robots while maintaining their stability and robustness, given their physical limitations?
In this paper, we introduce Advanced Expressive Whole-Body Control (ExBody2), an effective framework that maximizes the expressiveness of feasible whole-body motions on humanoid robots. The framework falls into a Sim2Real pipeline, where the policy will take a reference kinematic motion as input, and outputs the action that controls a real humanoid to conduct the motion in the real world. We train a single policy that generalizes across different input motions. We identify four technical designs to achieve this:
(i) Constructing feasible and diverse training datasets. Human motion datasets like AMASS [45] often contain complex movements beyond a robot’s physical capabilities, making tracking overly challenging and reducing performance. Some works address this by refining datasets. ExBody[8], for instance, filters out infeasible motions using language labels, though ambiguous descriptions (e.g., “dance”) can still include unsuitable movements. Other approaches, such as H2O[24] and OmniH2O [24], employ the SMPL model to simulate virtual humanoids and filter out complex motions. However, SMPL avatars can perform extreme actions that real robots cannot, creating a gap between simulation and real-world feasibility that can still impact training effectiveness. We systematically analyze dataset difficulty and diversity, focusing on upper-body movement diversity for stability and tracking accuracy, and lower-body feasibility for robust training. Experiments demonstrate that diversity and feasibility in datasets are crucial for expressive and stable behaviors in both training and testing phases.
(ii) Two-stage training. ExBody2 adopts a two-stage teacher-student training framework as illustrated in Figure 2. We first train a teacher policy using a standard RL algorithm, PPO [60] with privileged information, which includes real root velocity, accurate positions for each body link, and other physical properties (e.g. frictions) in simulation. This results in an effective policy to accurately mimic versatile human motion. We then learn a deployable student policy by performing DAgger[59]-style distillation on the teacher policy. For the student policy, the privileged information is replaced by a series of historical observations.
@@ 1 Introduction
(iii) Local keypoint tracking strategy. Previous whole-body tracking approaches, such as H2O [24] and OmniH2O [23], rely on global tracking of keypoint positions. This approach often leads to tracking failures in immediate next steps when robots struggle to align with current global keypoints, limiting their applications to highly stationary scenarios. In contrast, ExBody2 converts keypoints in the local frame, and decouples keypoint tracking from velocity control. To further improve the robustness of tracking, we periodically reset the global tracking keypoints to the local frame. This approach ensures sustained and robust whole-body tracking.
(iv) Long-range motion with CVAE. With enhanced whole-body tracking capabilities, our robot can now perform long-duration motion imitation. However, the motions in existing datasets are relatively short, presenting a challenge when aiming to enable continuous motion execution without manual resets. To overcome this, we train a Conditional Variational Autoencoder (CVAE) [64, 79] model to synthesize future motion. As shown in Figure 2, the CVAE takes on past motion information and generates future motion sequences, allowing the robot to execute complex and expressive movements seamlessly during deployment.
In the experiments, we evaluate four baseline methods on two robot platforms (Unitree G1 and H1) to assess the effects of different tracking methods, motion control strategies, and training techniques. Our approach demonstrates superior performance across full-body tracking accuracy, upper and lower body tracking, and velocity tracking, particularly excelling in velocity tracking and overall precision. Ablation studies further highlight that a curated dataset with moderate diversity and feasibility achieves optimal tracking performance, avoiding the drawbacks of overly simplistic or noisy datasets. By incorporating a teacher-student structure and an optimal history length, our approach achieves stable and efficient tracking in dynamic environments.
2 Related Work
@@
Humanoid Whole-Body Control Whole-body control for humanoid robots remains a challenging problem due to the system’s high non-linearity. Traditional approaches predominantly rely on dynamics modeling and control [49, 76, 28, 50, 12, 33, 73, 35, 26, 10, 11, 13, 51, 58]. Recent advances in reinforcement learning and sim-to-real transfer have demonstrated promising results in enabling complex whole-body skills for both quadrupedal and humanoid robots [48, 37, 18, 17, 16, 39, 63, 1, 75, 47, 14, 80, 7, 52, 72, 21, 74, 41, 42, 57, 19, 29, 6, 30, 61, 32, 31, 56, 66, 40, 62]. Notably, works such as [8, 24, 20] have focused on humanoid whole-body control for expressive motions, with applications in manipulation and imitation learning. However, these approaches still exhibit limitations in expressiveness and maneuverability, highlighting the untapped potential of humanoid robots.
Robot Motion Imitation Robot motion imitation can be categorized into two primary areas: manipulation and expressiveness. For manipulation tasks, robots—often wheeled or tabletop—prioritize precise control over balancing and ground contact, making humanoid morphology unnecessary. Such robots typically utilize data from teleoperation [79, 2, 3] or human demonstrations [70, 5, 65, 34]. In contrast, expressive motion imitation focuses on learning life-like behaviors from human or animal motion capture data. This task is more challenging due to the need for fine-grained control of contacts and balance. While reinforcement learning has enabled physics-based character motion imitation in simulation [54, 55, 67, 22, 44, 43, 77, 71, 68], transferring such methods to real robots remains a significant challenge [8, 23, 20, 25, 53, 16, 15].
Human Motion Data Human motion capture datasets [45, 4, 36] offer a rich source of reference motions for training low-level policies. Additionally, recent advances in generative modeling enable the creation of multi-modal motion data using diffusion models conditioned on text inputs [69, 78] and motion variational autoencoders (VAEs) [43]. Our work leverages these kinematically feasible motion datasets to train physically plausible controllers, while the acquisition of motion data remains outside our scope.
Figure 2: ExBody2’s framework. (a) ExBody2 presents a curated high-quality motion dataset, with highly diverse and feasible motions. (b) ExBody2 adopts a teacher-student learning framework on Unitree G1 and H1 whole-body control tasks. (c) ExBody2 employs a transformer-based CVAE for continuous motion synthesis. (d) By tracking the generated motion from CVAE, ExBody2 can be deployed seamlessly in the real world to mimic sustained motions expressively.
3 ExBody2: Learning Expressive Humanoid Whole-Body Control
We propose Advanced Expressive Whole-Body Control (ExBody2), a simple and effective sim-to-real framework for expressive and robust whole-body control. As illustrated in Figure 2, ExBody2 consists of four main components: dataset curation, policy learning, motion synthesis, and real-world deployment. In this section, we will introduce these components of ExBody2 in detail.
3.1 Motion Dataset Curation
In the curation of our motion datasets, significant emphasis is placed on the detailed analysis and selection of actions, particularly distinguishing between the capabilities required for upper versus lower body movements. This strategic approach aims to optimize the diversity of movements within feasible limits for robotic implementations.
Upper Body Actions Our dataset includes a wide array of upper body movements. This variety is crucial to challenge and enhance the robot’s adaptability and readiness for real-world applications, where flexibility in response to unexpected scenarios is vital.
Lower Body Actions The selection of lower body movements is approached with greater conservatism due to the mechanical limitations and stability requirements of robots. Basic locomotive actions such as walking and subtle stance adjustments are included. However, highly dynamic movements such as jumping or complex rotational movements are carefully assessed. Training on such infeasible actions would introduce noise into the learning process, thus reducing its effectiveness.
Balancing Act and Dataset Integrity The process of curating the datasets involves a delicate balance between challenging the robotic capabilities and maintaining actions within a feasible spectrum. Overly simplistic tasks could limit the training policy’s ability to generalize to new situations, while overly complex tasks might exceed the robot’s operational capabilities, leading to ineffective learning outcomes. Part of our dataset preparation, therefore, includes the exclusion or modification of entries that featured complex lower body movements beyond the robot’s capabilities. This step is crucial to ensure the datasets are rich in actionable content without overwhelming the robot’s learning algorithms with impractical tasks.
Through meticulous planning, we curate distinct datasets that are optimally balanced for robotic training. Extensive experimental validation will demonstrate the superiority of our dataset selection approach in Section 4.
3.2 Policy Learning
@@
ExBody2 aims at tracking a target motion more expressively in the whole body. To this end, ExBody2 adopts an effective two-stage teacher-student training procedure as in [38, 37]. Specifically, the oracle teacher policy is first trained with an off-the-shelf reinforcement learning (RL) algorithm, PPO [60], with privileged information that can be obtained only in simulators. For the second stage, we replace the privileged information with observations which are aligned with the real world, and distill the teacher policy to a deployable student policy. We train our policies using IsaacGym [46] with efficient parallel simulation.
3.2.1 Teacher Policy Training
@@
Privileged Information The privileged information $p_{t}$ contains some ground-truth states of the humanoid robot and the environment, which can only be observed in simulators. It contains the ground-truth root velocity, real body links’ positions, and physical properties (e.g. friction coefficients, motor strength). Privileged information can significantly improve the sample efficiency of RL algorithms, which is often leveraged to obtain a high-performing teacher policy.
Motion Tracking Target Similar to Exbody [9], ExBody2 learns a policy that can be controlled by the joystick commands (e.g. the linear velocity and body pose) when accurately tracking a whole-body motion. The motion tracking target consists of two components, which are (1) the desired joints and 3D keypoints in both the upper and lower body and (2) target root velocity and root pose. For the full information about the privileged information, motion tracking information, and proprioceptive observations for the teacher policy, please refer to the supplementary materials.
Reward Design Our reward function is meticulously constructed to enhance the performance and realism of the humanoid robot’s motion. The primary components of the reward include tracking the velocity, direction, and orientation of the root, alongside precise tracking of keypoints and joint positions. Additionally, we incorporate several regularization terms designed to boost the robot’s stability and enhance the transferability from simulation to real-world applications. The main elements of our tracking reward are detailed in Table 1, while supplementary rewards aimed at improving stability and sim2real capabilities would discussed extensively in the supplementary materials.
3.2.2 Student Policy Training

### v2

@@ Abstract
This paper tackles the challenge of enabling real-world humanoid robots to perform expressive and dynamic whole-body motions while maintaining overall stability and robustness. We propose Advanced Expressive Whole-Body Control (Exbody2), a method for producing whole-body tracking controllers that are trained on both human motion capture and simulated data and then transferred to the real world. We introduce a technique for decoupling the velocity tracking of the entire body from tracking body landmarks. We use a teacher policy to produce intermediate data that better conforms to the robot’s kinematics and to automatically filter away infeasible whole-body motions. This two-step approach enabled us to produce a student policy that can be deployed on the robot that can walk, crouch, and dance. We also provide insight into the trade-off between versatility and the tracking performance on specific motions. We observed significant improvement of tracking performance after fine-tuning on a small amount of data, at the expense of the others.
@@
Figure 1: Humanoid robot executing various expressive whole-body motions in the real world. The robot can (a) walk with a large stride from static standing, (b) dance along a long horizon (43 seconds) choreography, (c) dynamic sidestep with fluid weight shifts, (d) punch with different height configurations, (e) express various upper-body movements while maintaining balance, (f) powerful rightward body hook with dynamic shifts.
I Introduction
The premise of humanoid robots is to enable human-like motions while occupying human living spaces. However, a humanoid robot with human-level expressiveness and versatility that is also robust in maintaining stability and control remains elusive. Inherent to this problem is the dynamic and kinematic gap between robots and biological body structures and the need for the controller to make a trade-off between expressiveness and stability. How to let robots imitate human whole-body motion across this gap and achieve both is a key challenge.
This paper introduces Advanced Expressive Whole-Body Control (Exbody2), a framework that enables humanoid robots to perform expressive, human-like full-body motions with grace. At its core, Exbody2 features both a generalist and a specialist policy. The generalist policy, trained on diverse motion datasets, outperforms previous approaches by achieving high adaptability across a wide range of motions with a single policy. Building on this, we further fine-tune the policy for specific motion groups, producing specialist policies that ensure even higher fidelity in targeted behaviors. Our framework also incorporates a decoupled motion-velocity control strategy, allowing for both precise key body tracking and stable motion generation. These innovations set Exbody2 apart from previous methods and push the boundaries of humanoid motion control. Overall, there are three key innovations that distinguish Exbody2 from previous approaches:
(i) Generalist policy with automated data curation. Human motion datasets often contain movements beyond a robot’s physical limits, making tracking difficult and reducing performance. Some methods refine datasets, like ExBody [3] filtering motions via language labels, though ambiguous terms (e.g., “dance”) may still include infeasible actions. Others [18, 17] use SMPL avatars to simulate motions, but these can exceed real robot capabilities, impacting training. We identify the trade-off between dataset feasibility and diversity and develop an automated curation method that removes unsuitable lower-body motions while preserving diversity, enabling the policy to learn broad, expressive behaviors. Experiments validate that our method optimally balances feasibility and diversity, leading to improved stability and accuracy across diverse motion tasks.
(ii) Specialist policy with finetuning for targeted motions. While the generalist policy enables broad motion coverage, finetuning enhances precision for specific motion groups. Motions with similar patterns are easier to learn under a shared policy, as they require consistent control strategies and constraints. Instead of training from scratch, we refine the generalist policy, leveraging its learned priors for efficient adaptation. This allows the policy to better capture fine-grained motion details and improve tracking accuracy for specialized tasks. Additionally, motion labels or an action recognition model can classify input motions, enabling dynamic selection of the most suitable specialist policy.
@@ I Introduction
(iii) Decoupled motion-velocity control strategy. Previous whole-body tracking approaches, such as H2O [18] and OmniH2O [17], rely on global tracking of keypoint positions. This often leads to tracking failures in immediate next steps when robots struggle to align with current global keypoints, limiting their applications to highly stationary scenarios. In contrast, Exbody2 converts keypoints into the local frame and decouples keypoint tracking from velocity control. Velocity-based tracking guides movement, while key body tracking focuses on motion imitation, emphasizing expressive motion reproduction. To improve tracking robustness, Exbody2 adopts a teacher-student framework, where a teacher policy is first trained using PPO [51] with privileged information, including real root velocity, accurate body link positions, and physical properties (e.g., friction). The student policy is then trained via DAgger [50]-style distillation, using historical observations to infer privileged information, enabling real-world deployment without direct access to such data.
We evaluate our approach on the Unitree G1 against four state-of-the-art baselines, demonstrating higher fidelity across both simulation and real-world tests. In particular, the automated data curation yields a robust generalist policy that outperforms all previous methods as a single policy capable of handling diverse motions. Fine-tuning for specific tasks (e.g., dancing) further enhances motion quality and expressiveness. Overall, our results highlight ExBody2’s potential for bridging the gap between human-level expressiveness and reliable whole-body control in humanoid robots.
II Related Work
@@
Humanoid Whole-Body Control. Whole-body control for humanoid robots remains a complex and challenging problem due to the system’s high non-linearity and degrees of freedom. Traditional approaches predominantly rely on dynamics modeling and control [40, 62, 22, 41, 8, 25, 61, 27, 20, 6, 7, 9, 42, 49], which often require accurate system identification and physical modeling, and intensive online computation for real-time control to handle different external perturbations for locomotion stability. Recent advances in reinforcement learning (RL) and sim-to-real transfer have demonstrated promising results in enabling complex whole-body skills for humanoid robots [32, 33, 53, 10, 34, 35, 48, 23, 24, 47, 55, 52]. These approaches typically rely on training RL policies in simulation using task-specific rewards and environment randomization before transferring them to the real world. Notably, recent works such as [3, 18, 13] have advanced real-world humanoid whole-body control for expressive motion by incorporating human motion datasets [38] to guide RL training, with real-world applications such as motion imitation. However, these approaches still exhibit limitations in expressiveness and maneuverability, highlighting the untapped potential of humanoid robots. In contrast, our method enables the learning of more expressive and dynamic motions, enhancing the robot’s ability to perform complex whole-body movements.
Robot Motion Imitation. Robot motion imitation can be broadly categorized into manipulation and locomotion areas. For manipulation tasks, robots are often wheeled or tabletop, prioritizing precise control over balancing and ground contact, making humanoid morphology unnecessary. Such robots can imitate the motion through direct teleoperation [64, 14, 46], portable devices [59, 2, 5, 15] and learn from human videos with hand tracking or motion retargeting [58, 54, 31, 26]. In contrast, motion imitation for locomotion primarily aims to learn lifelike, natural behaviors from human or animal motion capture data. It requires precise control over contact dynamics, balance, and coordination across multiple degrees of freedom to achieve stable and realistic movement. While prior methods have enabled physics-based character motion imitation in simulation [44, 45, 56, 16, 37, 36, 63, 60, 57], transferring diverse motions to real robots [3, 17, 13, 19, 43, 12, 11] remains a significant challenge due to the hardware constraints. Previous methods [3, 17, 13, 19] typically rely on manually filtering feasible motion data with human effort or hand-crafted heuristics. However, manually filtered datasets may still contain infeasible motions or lack diversity, limiting the robot’s ability to fully utilize its hardware potential. Our method overcomes this challenge by automatically curating a diverse and feasible motion dataset, enabling more effective real-world deployment.
Figure 2: Exbody2’s framework. (a) Motion retargeting adapts raw human motion datasets to fit the humanoid robot’s morphology, generating a diverse set of training samples. (b) Automated dataset filtering ranks motions based on tracking errors and selects an optimal subset to train a generalist policy, balancing feasibility and diversity. (c) Specialist policy finetuning refines the generalist model for specific motion categories, such as walking, dancing, and kungfu, improving precision for targeted tasks. (d) The trained policies are deployed on a real humanoid robot, demonstrating expressive, dynamic, and stable whole-body motions in real-world environments.
III Exbody2: Learning Expressive Humanoid Whole-Body Control
We propose Advanced Expressive Whole-Body Control (Exbody2), a sim-to-real framework for expressive and robust whole-body control. As shown in Figure 2, Exbody2 first retargets human motion data to fit the robot’s morphology, then trains a generalist policy using an automated dataset curation strategy to balance feasibility and diversity. To improve precision on specific motion groups, we further fine-tune specialist policies and deploy it onto real humanoid robots. In the following sections, we detail our generalist-specialist training pipeline, and our policy structure design, the two main contributions of our work.
III-A Data-driven Generalist-specialist Training Pipeline
We adopt a Generalist–Specialist pipeline to balance adaptability and precision in whole-body motion tracking. At the heart of this pipeline is our Feasibility–Diversity Principle, which guides dataset design by requiring enough motion diversity (particularly in the upper body) to cover a broad distribution of tasks, while maintaining feasibility in the lower body to avoid unachievable or overly dynamic motions that degrade training stability. In practice, this amounts to filtering out extreme lower-body samples and retaining a wide range of upper-body actions. Further details on how we discover and validate this principle are in the appendix Section 3.
Generalist Policy. Leveraging the Feasibility–Diversity Principle, we train a generalist policy on a broad dataset that has been automatically pruned to remove infeasible lower-body motions. This ensures that the policy learns expressiveness in upper-body movements without being hindered by physically impractical transitions.
Specialist Policies. Once trained, the generalist policy can be fine-tuned for targeted tasks (e.g., dancing or kung fu), resulting in specialist policies with higher precision.
This two-step approach balances overall adaptability with focused accuracy, as we demonstrate in both simulation and real-world evaluations.
III-A1 Generalist policy with automated data curation
To obtain a policy $\pi$ that performs well across diverse motion inputs, we first train an initial policy $\pi_{0}$ on a comprehensive, unfiltered motion dataset $\mathcal{D}$, which is highly diverse with a lot of infeasible motions. After training $\pi_{0}$, we evaluate its tracking accuracy for each motion sequence $s\in\mathcal{D}$, obtaining a tracking error metric $e(s)$ that focuses on the lower body. The lower body plays a central role in dynamic feasibility and balance; thus, we focus on its tracking error for filtering. This preserves the upper body’s diversity while excluding excessively unstable motions, aligning with our feasibility–diversity principle. Specifically, we define
$e(s)\;=\;\alpha\,E_{\text{key}}(s)\;+\;\beta\,E_{\text{dof}}(s),$
where $E_{\text{key}}(s)$ is the mean keybody position error for the lower body (preventing extreme deviations such as flipping or rolling), and $E_{\text{dof}}(s)$ measures the mean joint-angle tracking error. The coefficients $\alpha$ and $\beta$ weight these two terms according to their relative importance for lower-body stability and precision. Once $e(s)$ is computed for each sequence, we rank the motions by their tracking errors and derive the empirical distribution $P(e)$.
The objective is to determine an error threshold $\tau$ such that the subset of motion sequences with $e(s)\leq\tau$, denoted as $\mathcal{D}_{\tau}=\{s\in\mathcal{D}\mid e(s)\leq\tau\}$, enables the training of a new policy $\pi_{\tau}$ that maximizes performance across the full dataset $\mathcal{D}$. Formally, we seek:
$\tau^{*}=\arg\max_{\tau}\mathbb{E}_{s\in\mathcal{D}}[\text{Performance}(\pi_{\tau},s)],$
where $\pi_{\tau}$ is trained on $\mathcal{D}_{\tau}$.
In practice, we divide $P(e)$ into evenly spaced error intervals to systematically evaluate the performance of policies trained on subsets corresponding to different thresholds $\tau$. Although we use a greedy search to identify the optimal threshold $\tau^{*}$, subsequent experiments reveal a strong trend in how the policy’s performance changes with $\tau$. When $\tau$ is too small, the filtered motions are overly simple, limiting the policy’s ability to generalize across the full dataset. Conversely, when $\tau$ is too large, the inclusion of many infeasible motions introduces noise, degrading the training effectiveness. The best-performing policy is consistently obtained at a moderate $\tau$, balancing diversity and feasibility.
The optimal threshold $\tau^{*}$, identified through this process, exhibits generalizability and can be effectively applied to other motion datasets, ensuring robust training and improved performance.
III-A2 Specialist policy with finetuning for targeted motions
After obtaining the generalist policy $\pi_{\tau^{*}}$, which balances diversity and feasibility across a broad range of motions, we further refine it into a specialist policy tailored for specific, high-precision tasks. This finetuning process leverages the pretrained generalist policy rather than training a new policy from scratch, offering several advantages.
First, specialist policies require tracking a smaller set of motions with higher precision, making finetuning a more efficient approach. Training directly from scratch for specialized motions is often impractical, as complex actions may be too difficult to learn without a strong prior. In contrast, the generalist policy provides a warm start, enabling the policy to adapt more effectively to challenging motions.
Second, the generalist policy has been exposed to a wider range of motion sequences, improving robustness to variations and unexpected disturbances. While a policy trained from scratch on specific motions may perform well in simulation, it lacks exposure to diverse conditions. As a result, when deployed in the real world, such a policy may struggle with unseen disturbances or novel states, leading to failures. By finetuning from a generalist model, the specialist policy inherits adaptability and robustness, improving its real-world generalization.
Finally, finetuning can reduce training time and computational requirements, as it refines an already well-trained model instead of learning from scratch. This makes it a practical and scalable approach for developing high-precision policies for specialized tasks while maintaining strong generalization capabilities.
Figure 3: Teacher-student framework for humanoid motion learning, where the teacher uses privileged information, and the student learns from past observations to generate control actions.
III-B Policy Objective and Architecture
@@
Exbody2 aims at tracking a target motion more expressively in the whole body. To this end, Exbody2 adopts a two-stage teacher-student training procedure as in [29, 28]. Specifically, the oracle teacher policy is first trained with an off-the-shelf reinforcement learning (RL) algorithm, PPO [51], with privileged information that can be obtained only in simulators. For the second stage, we replace the privileged information with observations which are aligned with the real world, and distill the teacher policy to a deployable student policy. We train our policies using IsaacGym [39] with efficient parallel simulation.
III-B1 Teacher Policy Training
@@
Privileged Information. The privileged information $p_{t}$ contains some ground-truth states of the humanoid robot and the environment, which can only be observed in simulators. It contains the ground-truth root velocity, real body links’ positions, and physical properties (e.g. friction coefficients, motor strength). Privileged information can significantly improve the sample efficiency of RL algorithms, which is often leveraged to obtain a high-performing teacher policy.
Motion Tracking Target. Similar to Exbody [4], Exbody2 learns a policy that can be controlled by the joystick commands (e.g. the linear velocity and body pose) when accurately tracking a whole-body motion. The motion tracking target consists of two components, which are (1) the desired joints and 3D keypoints in both the upper and lower body and (2) target root velocity and root pose. For the full information about the privileged information, motion tracking information, and proprioceptive observations for the teacher policy, please refer to the supplementary materials.
Reward Design. Our reward function is carefully constructed to enhance the performance and realism of the humanoid robot’s motion. The primary components of the reward include tracking the velocity, direction, and orientation of the root, alongside precise tracking of keypoints and joint positions. Additionally, we incorporate several regularization terms designed to boost the robot’s stability and enhance the transferability from simulation to real-world applications. The main elements of our tracking reward are detailed in Table I, while supplementary rewards aimed at improving stability and sim2real capabilities would discussed in the supplementary materials.
III-B2 Student Policy Training